OpenStudy (anonymous):

Solve by completeing the square: 8x^2 - 80x = 32. Can you help me work this out? I have the answer, just not sure how to solve it with all the steps.

OpenStudy (anonymous):

8x^2 - 80x = 32 Divide both sides by x^2 - 10x = 4 x^2 - 10x + 25 - 25 = 4 (x-5)^2 -25 + 25 = 4+25 (x-5)^2 = 29 \(x-5 = \pm \sqrt{29}\) \(x = 5\pm \sqrt{29}\)

OpenStudy (anonymous):

x^2-10x=4 To complete the square we (b/2)^2 and in this case b =10 need to add 25 that is 10/2=5^2 x^2-10x+25=4+25 (x-5)^2=29 \[x-5= \pm \sqrt{29}\] \[x=5 \pm \sqrt{29}\]

OpenStudy (anonymous):

 I GOT ALSO THE SAME ANSWER..